OpenStudy (anonymous):

D=1.2sqr(h) Solve equation for h.

myininaya (myininaya):

divide 1.2 on both sides then square both sides

OpenStudy (anonymous):

correct..that is D/1.2=sqrth..squaring both sides gives D^2 /(1.2)^2=h D^2 /1.44 = h answer